INAF LBC MODS queue observing, 2013 January 19-20

Observer: M.Faccini, F.Cusano
Telescope Operator: G.Bechetti
Support Astronomer: O. Khun
Instrument: LBC + MODS

Summary

Clear sky, seeing 0.8 arcsec, moon at the first quarter. We started with LBC. Problem at the beginning of the night, stars elongated in the blue channel, we tried to fix it manually (more than one hour), but it did not work. We switch to MODS.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Annibali 23B hd74377 Dual grating 2x1s Lick standard
Annibali 23B hd84937 Dual grating 1s Lick standard
spectrophotometric standard GD71 Dual grating 3x180s  
spectrophotometric standard Hiltner 600 Dual grating 3x60s  
spectrophotometric standard G191-B2B Dual grating 3x90s  
Annibali 23B NGC4449 Mos dual grating 2h science target
Fritz 4A LensGG243 Mos dual grating 2h science target COMPLETED
Annibali 23B Feige66 Dual grating 1x120s  
Gilli 18D Feige66 Red grating 3x90s  
Gilli 18D cxoj12 Red grating 4x900s science target;file 0058 wrong filter; 0059 correct filter but header wrong
Annibali 23B hd160693 Dual grating 1s Lick standard

Details

01:32 START WITH LBC

01:33 slew to CalchiNovati 3A, M31

01:35 dopfia first X2

01:45 copointing

01:55 dopfia X2

02:04 dopfia

02:08 problem with dopfia, sending again

02:17 launched M31_Lensing_60s_1.ob

02:27 ob stopped, elondated stars in the blue channel, we try dopfia again on the pointing ob M31_Focus.ob

02:31 we are trying to send manually corrections to fix the elongation

02:35 dopfia

02:42 test ob, stars still elongated, images b024302 and r024303

02:47 sending manual correction

03:06 dopfia

03:28 problem not solved

03:20 SWITCH TO MODS

04:00 slew to Annibali 23 B

04:00 standard star, acqMods hd74377_lick_runb.acq mods1r.20130120.0004.fits

04:07 modsAlign -l mods1r.20130120.0003.fits mods1r.20130120.0004.fits

04:14 execMods hd74377_lick_runb.obs mods1r.20130120.0005.fits/mods1r.20130120.0006.fits, mods1b.20130120.0002.fits/mods1b.20130120.0003.fits

04:37 slew to standard star GD71

04:41 acqMods GD71.acq mods1r.20130120.0009/0011

04:50 execMods GD71.obs mods1r.20130120.0012/0015, mods1b.20130120.0004/0006, the script took one spectra more in the red

05:13 acqMods hd84937_lick_runb.acq mods1r.20130120.0016/0021

05:22 modsAlign -l mods1r.20130120.0016.fits mods1r.20130120.0017.fits, did not work propertly, we gave some corrections by hand

05:33 execMods hd84937_lick_runb.obs mods1r.20130120.0022, mods1b.20130120.0007

05:47 slew to standard Hiltner 600

05:48 acqMods hiltner600.acq mods1r.20130120.0023/0024

06:02 execMods hiltner600.obs mods1r.20130120.0025/0027, mods1b.20130120.0008/0010

06:18 stars elongated in the upper right and lower left of MODS images (problem recognized during the acquisitions)

06:29 slew to standard star G191-B2B

06:30 acqMods g191b2b.acq mods1r.20130120.0030/0031

06:39 execMods g191b2b.obs mods1r.20130120.0032/0034, mods1b.20130120.0013/0015

06:59 slew to target NGC4449 Annibali 23 B

06:59 acqMods mods.1.4449runb.acq mods1r.20130120.0036/0038

07:15 execMods mods.1.4449runb.obs mods1r.20130120.0039/0042, mods1b.20130120.0017/0020

09:25 slew to target Fritz 4A

09:25 acqMods FritzLensGG243.acq mods1r.20130120.0043/45

09:41 execMods FritzLensGG243.obs mods1r.20130120.0046/47, mods1b.20130120.0021/0022

11:17 acqMods feige66.acq (standard both for Annibali 23B and Gilli 18D) mods1r.20130120.0048/50

11:28 execMods feige66.acq, dual grating, mods1r.20130120.0051, mods1b.20130120.0023

11:40 execMods feige66.acq, red grating, mods1r.20130120.0052/54

11:46 slew to Gilli 18D, acqMods cxoj12.acq mods1r.20130120.0055/0057

12:08 execMods cxoj12.obs mods1r.20130120.0058/0061 spectra 0058 has the wrong filter (filter used zSDSS), there was an error in putting the correct filter, error message Rfilter move fault; spectra 0059 in the header has filter zSDSS but it was acquired using the correct filter GC495

13:20 slew to Lick standard of Annibali, acqMods hd160693_lick_runb.acq mods1r.20130120.0062/0064

13:30 execMods hd160693_lick_runb.obs mods1r.20130120.0065, mods1b.20130120.0024

13:50 Starting calibrations -- MarcoFaccini - 19 Jan 2013
Topic revision: r9 - 29 Jan 2013, RobertoSpeziali
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback